### v1

# Beyond Poisson: First-Passage Asymptotics of Renewal Shot Noise

Abstract
The first-passage time (FPT) of a stochastic signal to a threshold is a fundamental observable across physics, biology, and finance. While renewal shot noise is a canonical model for such signals, analytical results for its FPT have remained confined to the Poisson (Markovian) case, despite the prevalence of non-Poisson arrival statistics in applications from neuronal spiking to gene expression. We break this long-standing barrier by deriving the first universal asymptotic formula for the mean FPT $\langle T_{b}\rangle$ to reach level $b$ for renewal shot noise with general arrival statistics and exponential marks. Our central result is a closed-form expression that reveals precisely how general inter-arrival statistics impact the naive Arrhenius law. We show that the short-time behavior of the interarrival distribution dictates universal scaling corrections, ranging from stretched-exponential to algebraic, that can dramatically accelerate threshold crossing. Furthermore, we argue and confirm numerically that the full FPT distribution becomes exponential at large thresholds, implying that $\langle T_{b}\rangle$ provides a complete asymptotic characterization. Our work, enabled by a novel exact solution for the moments of the noise, establishes a general framework for analyzing extreme events in non-Markovian systems with relaxation.
Threshold-crossing events driven by stochastic jump-decay processes are ubiquitous across physics, biology, and finance. In neurons, spikes occur when membrane voltage exceeds a threshold between synaptic inputs [1, 2, 3, 4, 5]; in gene expression, bursty mRNA/protein levels must cross regulatory thresholds to trigger phenotypic switching [6, 7, 8, 9, 10, 11, 12]; in materials science, stress fluctuations trigger yielding events with relaxation between avalanches [13, 14]; and in finance, barrier crossings determine option pricing and ruin probabilities [15, 16]. In all these contexts, the first-passage time (FPT) of the noise process $X(t)$ to a threshold $b$ is the central observable.
The natural model capturing these dynamics is renewal shot noise, defined by
$X(t)=\sum_{t_{i}\leq t}x_{i}\,e^{-\gamma(t-t_{i})},$ (1)
where $x_{i}$ are i.i.d. marks and interarrival times $\tau_{i}=t_{i+1}-t_{i}$ are i.i.d. with density $w(\tau)$ (see FIG. 1). This model embodies two essential features: impulsive bursts at random times, and relaxation between events. The classical Poisson case ($w(t)=re^{-rt}$) renders $X(t)$ Markovian, and its FPT statistics are well understood [17, 18, 19, 20, 21]. However, many applications exhibit strongly non-Poissonian arrival statistics, such as refractory periods in neuronal spiking or bursty transcription in gene expression [22, 23, 24, 25, 26, 11], which render the process genuinely non-Markovian. Despite decades of study, analytical progress on FPT statistics has remained confined to the Poisson case, with non-Poisson shot noise presenting a long-standing challenge, as is often the case for non-Markovian processes [27, 28, 29, 30]. While general, exact integral equations satisfied by the MFPT are known [21], these have proven to be intractable beyond the Poisson case.
Figure 1: A typical realization of renewal shot noise $X(t)$ (solid line). The vertical green line signals the FPT $T_{b}$, where $X(t)$ exceeds the threshold $b=5$ (horizontal red line) for the first time.
In this Letter, we break this barrier by deriving the first exact asymptotic expression for the mean FPT $\langle T_{b}\rangle$ for general renewal arrivals and exponential marks. Our result, given by a compact product formula Eq. (2), reveals explicitly how general interarrival statistics impact the FPT scaling, and in particular reveals universal deviations from the naive Arrhenius law. This breakthrough is enabled by a novel closed-form expression for the Laplace transform of all moments of $X(t)$, a result of independent interest that provides a powerful analytical tool for studying renewal shot noise.
Main result. We now present our main result for the MFPT $\langle T_{b}\rangle$. The marks are taken to be exponentially distributed, $x_{i}\sim\mathrm{Exp}(\lambda^{-1})$, and the arrival process has a finite mean rate $r\equiv\left(\int_{0}^{\infty}t\,w(t)\,dt\right)^{-1}$. The process may start at any value $0\leq X(0)\ll b$. For a function $f(t)$, we denote its Laplace transform by $\hat{f}(s)=\int_{0}^{\infty}e^{-st}f(t)\,dt$. Our central result is the following simple, exact asymptotic expression for the MFPT at large thresholds  [^1] The product term $\prod_{m=1}^{\lambda b}\!\bigl[1-\hat{w}(m\gamma)\bigr]$ is defined via analytic continuation when $\lambda b\notin\mathbb{N}$.:
$\langle T_{b}\rangle\;\sim\;\frac{\exp(\lambda b)}{r}\prod_{m=1}^{\lambda b}\!\bigl[1-\hat{w}(m\gamma)\bigr],\qquad b\to\infty.$ (2)
Figure 2: MFPT $\langle T_{b}\rangle$ of renewal shot noise with exponential marks (symbols: simulations; dashed lines: exact asymptotics, Eq. (2)). Interarrival times follow a Gamma distribution $w(t)=\tfrac{rk}{\Gamma(k)}(rkt)^{k-1}e^{-rkt}$ with mean rate $r$ and shape $k$, while marks are exponential with $\lambda=1$. Cases with $k=3$ (blue, $\gamma=1,r=0.4$) and $k=2$ (yellow, $\gamma=1.5,r=0.6$) highlight refractory effects ($w(0)=0$), whereas $k=0.75$ (green, $\gamma=4.5,r=1.6$) and the Poisson limit $k=1$ (red, $\gamma=2.5,r=1.2$) illustrate bursty dynamics. Statistical errors are smaller than the symbol size. The inset shows the simulated/theoretical ratio, confirming good quantitative agreement even at moderate $\lambda b$, with convergence speed set by $k$: larger $k$ suppresses short interarrivals and accelerates approach to the asymptotics.
This key formula, confirmed numerically in FIG. 2, leads to several important insights. (i) We check that Eq. (2) reduces to known results in important limits. First, for Poisson arrivals, $\hat{w}(s)=r/(s+r)$, it simplifies to the classical asymptotic expression [17, 18, 19, 20, 21]:
$\langle T_{b}^{\mathrm{Poisson}}\rangle\underset{b\to\infty}{\sim}\frac{1}{\gamma}\,\Gamma\!\left(\tfrac{r}{\gamma}\right)\,(\lambda b)^{-r/\gamma}\,e^{\lambda b}.$ (3)
Second, in the limit of instantaneous relaxation ($\gamma\to\infty$), we have $\hat{w}(m\gamma)\to 0$, yielding the pure Arrhenius law $r\langle T_{b}\rangle\sim e^{\lambda b}=1/\mathbb{P}(x>b)$. Indeed, in this limit, each impulse is an upcrossing with probability $\mathbb{P}(x>b)=e^{-\lambda b}$. (ii) Equation (2) extends MFPT asymptotics for shot noise far beyond the Poisson case, providing (to our knowledge) the first closed analytic form valid for general renewal arrivals and exponential marks. Comparable asymptotics are extremely rare for non-Markovian processes [30, 29, 27, 32]. (iii) Although the full distribution of $T_{b}$ for arbitrary $b$ remains an open question for non-Markovian shot noise, our results provide a complete asymptotic characterization for large values of $b$, where we argue that crossings become asymptotically independent, implying that $T_{b}$ follows an exponential distribution. While this has been rigorously established only for Poisson arrivals [17, 18], it is a common phenomenon in rare-event limits [33, 34]. Our simulations FIG. 3 confirm that for large $b$,
$\mathbb{P}(T_{b}>t)\;\underset{b\to\infty}{\sim}\;\exp\!\left(-t/\langle T_{b}\rangle\right),$ (4)
where $\langle T_{b}\rangle$ is now fully explicit (2). Thus, under the physically sound hypothesis that upcrossings become independent, our result (2) fully quantifies FPT statistics of shot noise with exponential marks in the large-threshold limit.
Figure 3: Numerical confirmation of the exponential distribution of $T_{b}$ for large $b$. The cumulative distribution function $\mathbb{P}(T_{b}<t)$ from simulations (in blue) is shown for a threshold $b=8$, with Gamma-distributed interarrivals (shape $k=2$, rate $r=0.6$), decay $\gamma=1.5$, and mark rate $\lambda=1$. The black dotted line is the expected exponential distribution $1-\exp(-t/\langle T_{b}\rangle)$, with $\langle T_{b}\rangle$ given by Eq. (2). The excellent agreement strongly supports our assertion (4) that the FPT distribution is asymptotically exponential for large $b$.
(iv) A central question in first-passage phenomena is how the MFPT deviates from the simple Arrhenius law $\langle T_{b}\rangle\propto e^{\lambda b}$ [30, 35, 36]. Assume the interarrival density behaves for short times as $w(t)\sim c\,t^{\kappa-1}$ with $\kappa>1/2$ (the case $0<\kappa\leq 1/2$ yields more intricate, stretched-exponential expressions and is treated in the SM). Our expression (2) then yields the universal asymptotic scaling
$\langle T_{b}\rangle\underset{b\to\infty}{\sim}\frac{B\,e^{\lambda b}}{r}\times\begin{cases}e^{-\tfrac{c\,\Gamma(\kappa)}{(1-\kappa)\,\gamma^{\kappa}}\left(\lambda b\right)^{1-\kappa}},&\tfrac{1}{2}<\kappa<1,\\[4.0pt] (\lambda b)^{-c/\gamma},&\kappa=1,\\[4.0pt] 1,&\kappa>1,\end{cases}$ (5)
where the prefactor $B$ (given in the SM) is nonuniversal and depends on the full interarrival law. The scaling itself, however, is universal and dictated solely by the short-time behavior of $w(t)$ through the parameters $c$ and $\kappa$. Physically, when $w(0)=0$ ($\kappa>1$), short interarrival times are suppressed, akin to refractory periods in neuronal spiking. In this regime, exponential relaxation prevents the accumulation of impulses, and threshold crossings are dominated by a single, rare large mark of size $\mathcal{O}(b)$, resulting in pure Arrhenius scaling. Conversely, when short gaps are prevalent ($\kappa\leq 1$), as observed in bursty gene transcription, crossings are driven by bursts of arrivals rather than isolated events. This cooperative mechanism leads to significant reductions in the mean first-passage time relative to the Arrhenius form: algebraic corrections when $\kappa=1$, and stretched-exponential corrections when $\kappa<1$.
Derivation of (2). The derivation of our main result (2) is based on a new, exact formula for the moments of $X(t)$, a result of independent interest. Despite extensive study of shot-noise [37, 38, 39, 16], closed-form results for the moments of renewal shot noise remain limited. The Poisson case, where $X(t)$ is Markovian, is classical and well understood [40, 41], but for general arrivals, analysis has typically stopped at the first two moments [42, 41, 22]. Higher moments have previously been addressed through recursive schemes in actuarial mathematics [43, 39], with much of this line of analysis dating back to Takács [44], but such schemes become unwieldy at increasing order. In neuroscience, where shot noise corresponds to postsynaptic currents in the so-called Stein model [45, 2, 1] and non-Poisson arrival statistics are well established [22], analysis typically relies on Gaussian or diffusion approximations [2] which are known to fail outside narrow parameter regimes [2, 22]. Although more detailed pool-based synaptic release and network models have been analyzed in detail [46, 4], the renewal shot noise statistics treated here are more general and not contained in those models. Consequently, despite decades of use across disciplines, no compact closed-form expression for higher moments has been available.
We now present a closed-form expression for the Laplace transform of $\langle X(t)^{n}\rangle$, valid for arbitrary renewal processes and mark distributions with finite moments. Let $\mu_{k}\equiv\langle x_{i}^{k}\rangle$ and $\hat{\psi}(s)\equiv\hat{w}(s)/[1-\hat{w}(s)]$. We show in SM the following exact expression:
$\widehat{\langle X(t)^{n}\rangle}=\frac{1}{s+n\gamma}$ $\sum_{\begin{subarray}{c}n_{1}+\dots+n_{k}=n\\ 1\leq k\leq n\end{subarray}}\binom{n}{n_{1},\dots,n_{k}}\times$
$\prod_{m=1}^{k}\mu_{n_{m}}\cdot\hat{\psi}\left(s+\gamma\sum_{j=1}^{m-1}n_{j}\right).$ (6)
For the crucial case of exponential marks ($\mu_{m}=m!\,\lambda^{-m}$), this general result collapses to a remarkably simple product form:
$\widehat{\langle X(t)^{n}\rangle}=\frac{\hat{\psi}(s)\,\lambda^{-n}\,n!}{s+n\gamma}\prod_{m=1}^{n-1}\left(1+\hat{\psi}(s+m\gamma)\right).$ (7)
The structure of Eq. (Beyond Poisson: First-Passage Asymptotics of Renewal Shot Noise) reflects the underlying physics: it sums over all ways to distribute $n$ impulses among $k$ distinct arrivals ($1\leq k\leq n$). Each partition $n_{1}+\dots+n_{k}=n$ corresponds to a specific clustering pattern, with the $\hat{\psi}-$factors encoding the temporal structure.
To the best of our knowledge, neither (Beyond Poisson: First-Passage Asymptotics of Renewal Shot Noise) nor its specialization (7) has appeared in the literature. We recognize the latter as closely connected to the factorial moments of the $G/M/\infty$ queue [44, 12] (see SM), but it does not appear in recent reviews of shot noise [38]. These expressions offer a broadly applicable, fully analytical alternative to the recursive or approximate methods commonly used. In particular, the final-value theorem $\lim\limits_{s\to 0}s\hat{f}(s)=\lim\limits_{t\to\infty}f(t)$ applied to (Beyond Poisson: First-Passage Asymptotics of Renewal Shot Noise) gives exact moments of all orders in the stationary state $X(t\to\infty)\equiv X_{\infty}$ of the shot noise. As shown below, it is precisely these stationary moments that will allow us to compute the MFPT. Our starting point is the following rare-event estimate for the MFPT to threshold $b$:
$\langle T_{b}\rangle\;\mathrel{\mathop{\kern 0.0pt\sim}\limits_{b\to\infty}}\;\frac{1}{r\,p(b)},$ (8)
where $p(b)$ is the probability that a single impulse in the stationary state pushes the process above the threshold $b$. Equation (8) rests on three key points. (i) For large $b$, crossings are rare and thus take a long time, so the process is near stationarity when a crossing occurs. (ii) Multiple crossings are exponentially less likely than a single crossing in the large-$b$ limit, so a crossing of $b$ at some time $t$ is very likely to be the first one. (iii) In the stationary regime, all impulses are equivalent, so the mean number of crossings per unit time is rp(b)r\,p(b). The quantity of interest is now
$p(b)\equiv\mathbb{P}(X^{+}_{\infty}>b,\;X^{-}_{\infty}<b),$ (9)
where $X^{-}_{\infty}$ (resp. $X^{+}_{\infty}$) denotes the shot noise just before (resp. after) an impulse in the stationary regime. Importantly, except in the case of Poisson arrivals, $X^{-}_{\infty}$ is not distributed as $X_{\infty}$. Because the mark $X^{+}_{\infty}-X^{-}_{\infty}$ is exponentially distributed with mean $1/\lambda$, Eq. (9) can be written as
$p(b)=e^{-\lambda b}\,\int_{0}^{b}e^{\lambda x}\,P(X^{-}_{\infty}=x)\,dx,$ (10)
where we introduced the truncated moment–generating function (mgf) of the pre-burst shot noise. In the SM we show that a computation analogous to Eq. (Beyond Poisson: First-Passage Asymptotics of Renewal Shot Noise) gives the stationary pre-burst moments
$\langle(X^{-}_{\infty})^{n}\rangle=\hat{\psi}(n\gamma)\,\lambda^{-n}\,n!\,\prod_{m=1}^{n-1}\!\left(1+\hat{\psi}(m\gamma)\right),\qquad n\geq 1.$ (11)
The final step relies on an asymptotic duality (derived in SM) between truncated moment sums and integrals: for a random variable $Y$ with mgf finite for $t<\lambda$,
$\sum_{n=0}^{\lambda b}\frac{\langle(\lambda Y)^{n}\rangle}{n!}\;\underset{b\gg 1/\lambda}{\sim}\;\int_{0}^{b}e^{\lambda y}P(Y=y)\,dy.$ (12)
Applying (12) to $Y=X_{\infty}^{-}$ and noting the telescopic identity for $n\geq 1$,
$\frac{\langle(\lambda X^{-}_{\infty})^{n}\rangle}{n!}=B_{n+1}-B_{n},\quad B_{n}\equiv\prod_{m=1}^{n-1}[1-\hat{w}(m\gamma)]^{-1},$ (13)
the sum in (12) collapses, finally yielding Eq. (2).
Physical interpretation. We emphasize that the product form in Eq. (2) is not a mere mathematical artifact but reflects a clear physical mechanism, presented below. For large thresholds $b$, the MFPT is given by $\langle T_{b}\rangle\sim 1/(r\,p(b))$. To understand the product structure physically, we analyze how the crossing probability changes when we increment the threshold by one mark unit: $b^{+}\equiv b+1/\lambda$. The ratio $p(b^{+})/p(b)$ represents the conditional probability to cross $b^{+}$ given that $b$ has been crossed, and is governed by two distinct scenarios. Consider the overshoot $x_{0}$ remaining immediately after crossing $b$. Crucially, for exponential marks, $x_{0}$ is also exponentially distributed due to the memoryless property. This leads to two possible mechanisms: (S1) With probability $e^{-1}$, we have $x_{0}>1/\lambda$, meaning the same impulse that crossed $b$ also suffices to cross $b^{+}$. This contributes a term $e^{-1}p(b)$ and is responsible for the Arrhenius factor $e^{\lambda b}$. (S2) If instead $x_{0}\leq 1/\lambda$, the process starts below $b^{+}$ after crossing $b$. Since $X(t)$ rarely sits near such high values, crossing $b^{+}$ typically occurs through a burst of additional impulses before significant relaxation below $b$ can occur. This burst mechanism explains deviations from pure Arrhenius scaling. More precisely, in scenario (S2), crossing $b^{+}$ is achieved by $n\geq 1$ additional impulses with amplitudes $x_{1},\dots,x_{n}$ arriving at times $t_{1},\dots,t_{n}$. The interarrival times $\tau_{i}=t_{i}-t_{i-1}$ are most probably smaller than $(\lambda b^{+}\gamma)^{-1}$, the time needed for $X(t)$ starting from $b^{+}$ to relax by one mark unit $1/\lambda$. Since $\gamma\tau_{i}\ll 1$ for these relevant timescales, the condition for crossing $b^{+}$ after $n$ additional impulses becomes:
[…]
This condition ensures that the process remains between $b$ and $b^{+}$ until the final impulse pushes it above $b^{+}$. The total weight of scenario (S2) is therefore:
$\sum_{n=1}^{\infty}\int_{0}^{\infty}dt\int_{\tau_{1}+\cdots+\tau_{n}=t}w(\tau_{1})\cdots w(\tau_{n})\,\mathbb{P}(A_{n}).$ (15)
While this expression holds for general mark distributions, it simplifies dramatically for exponential marks. In this case,
$\mathbb{P}(A_{n})=e^{-\lambda\gamma b^{+}\sum_{k=1}^{n}\tau_{k}-1},$
which yields a term $\hat{w}(\lambda b^{+}\gamma)^{n}$ in (15). Combining both scenarios yields the recursion relation:
$\frac{p(b^{+})}{p(b)}\underset{b\to\infty}{\sim}\frac{1}{e}\left(1+\frac{\hat{w}(\lambda b^{+}\gamma)}{1-\hat{w}(\lambda b^{+}\gamma)}\right),$ (16)
which is exactly equivalent to our main result in Eq. (2)  [^2] Because only interarrival times with $\gamma\tau_{i}\ll 1$ matter, the MFPT expression (2) depends only on the local decay slope after an impulse, not on the full exponential form. The same asymptotics therefore hold in the large-$b$ limit if exponential relaxation is replaced by any smooth relaxation with the same initial slope. See [48] for the case of a constant slope.. This physical picture also clarifies why extending the explicit MFPT beyond the exponential–mark case solved here appears out of reach: for general marks the overshoot law is unknown and $\mathbb{P}(A_{n})$ lacks a closed form.